# Active particles in tunable crowded environments

### v1

@@ 2.2 Active particles in tunable colloidal environments
Figure 7: Reorientation in quenched environments leading to chiral motion. (a) Mean torque experienced by an active Janus particle swimming throughout a monolayer at area fraction $\phi=0.48$ for different applied electric fields. Solid and empty symbols correspond to experiments and simulations, respectively. The red dashed line connects the numerical data and the black solid line marks the linear relationship between $\rm\langle M\rangle$ and $\rm E^{2}$. (b) Persistence time $\tau$ as a function of the average velocity $v$ in colloidal environments at different $\phi$, as indicated in the legend. As in (a), filled symbols are obtained from experiments, while empty symbols are from numerical simulations matching the experimental values of $v$. (c) Helical active trajectory of 44 minutes duration observed in experiments when the area fraction of the surrounding colloidal bath is $\phi=0.64$ (see microscopy image). The scale bar corresponds to 2020 $\rm\mu m$. (d) Numerical trajectory of an active particles swimming at velocity $v=2.4$ $\rm\mu m/s$ in colloidal environments at $\phi=0.35$. The color indicates the time: from $\rm t=0$ (light blue) to $\rm t=40$ $\rm s$ (blue). The lever arm $\ell=0.4$ $\rm\mu m$ determining the torque is sketched as blue arrow pointing from the geometric particle center to the interaction center. The scale bar is 22 $\rm\mu m$ in (c) and 11 $\rm\mu m$ in (d).
@@ 2.3 Spontaneous chiralization
The scenario depicted so far also occurs over a broad range of packing fractions $\phi$. Figure 7(b) summarizes the dependence of $\tau$ on $v$ in colloidal environments at different packing fraction: $\rm\phi=0.15$ (green data), $\rm\phi=0.35$ (blue data) and $\rm\phi=0.48$ (red data, redrawn from Fig. 5(e)). In all instances, we report a drop of persistence time as the swimming velocity increases. Only for small $v$ and $\phi$ (first three green data points) or large $v$ and $\phi$ (last two red data points), $\tau$ remains roughly constant. The observed behavior is reproduced by numerical simulations (empty symbols and connecting dotted lines) remarkably well. Note, however, that the simple model described in Fig. 6 breaks down if the environment is too sparse (at $\phi=0.15$ for small $\rm K$ values) and for strong velocity fluctuations, i.e. upon caging (at $\phi=0.48$ and $\phi=0.64$, for large $\rm K$ values). Finally, at $\phi=0.64$ a significant number of active particles undergo helical motion (Fig. 7(c) and Supplementary Video S7). They behave as if they were chiral, with orbit’s radii as large as $\sim 5\rm R_{p}$, but no preferential kick-off direction. Because both the translational and angular velocity grow linearly with $\rm E^{2}$ (Fig. 4(a) and Fig. 7(a)), the radius of the orbit, $v/\langle\dot{\theta}\rangle$, does not depend on $\rm E$. The same break of polar symmetry is observed in numerical simulations but, at $\phi=0.64$, the spontaneous chiralization is hindered by strong caging effects which, as discussed above, are more pronounced in simulations than experiments. To recover the same phenomenology reported in experiments, we keep the packing fraction relatively low ($\phi=0.35$) and virtually shift the point on which dipolar interactions act from the center of the $\rm SiO_{2}$ hemisphere towards the edge (see blue arrows in the sketch of Fig. 7(d)). Practically, this operation increases the lever arm $\ell$ and amplifies the torque exerted by the passive particles due to dipolar interactions, while keeping the swimming velocity and stiffness of the medium fixed. As the torque increases, the active colloid is bound to make a helical motion – a signature of chirality. As in experiments, there is no preferential direction of rotation.
@@
In agreement with the experimental and numerical results, $\langle\dot{\theta}\rangle$ linearly increases with the self-propulsion speed $v_{0}$, and thus vanishes in equilibrium or for a force that is applied to the geometric center of the particle, such that $\ell=0$. In addition, this term disappears when the local density change of the environment around the active particle is negligible, as in experiments with a low-density background. The full coarse-grained theory is derived and reported in Sec. 4.
@@ 3 Discussion and Conclusions
After presenting an experimental strategy to tune the structure of colloidal monolayers at fixed packing fraction, we investigated the dynamical properties of active colloids cruising throughout them. The active motion is strongly affected by the environment, which alters the particle speed and reorientation, and even gives rise to unexpected helical motion. The mechanical properties of the environment and microswimmers activity are adjusted reversibly and in situ using the same experimental control knob: the magnitude of an applied AC electric field.
@@
The role of orthogonal AC electric fields in determining the pair interaction between charged microparticles sitting onto planar electrodes has been extensively investigated in the past decades [53, 54, 55, 56]. In particular, the AC frequency is of paramount importance since it regulates whether the interactions between particle pairs are attractive [57, 58, 59, 60] or repulsive [56, 43, 61]. At our operating frequency $\rm f=20$ $\rm kHz$, long-range dipole-dipole repulsive interactions between $\rm SiO_{2}$ particles are dominant; their strength ($\rm K$, Eq. (9)) is always much larger than the energy scale of the WCA potential ($\epsilon$, Eq. (8)) at the typical experimental inter-particle distances identified by the first peak of the pair correlation function (Fig. 2(a), Fig. S1 and Table S1). As such, the experimental system presented in Figs. 2-3 is similar to the colloidal monolayers described in Refs. [56, 62, 63], where quenching (or melting) took place under external electric ($\rm f>1$ $\rm MHz$) and magnetic fields. These works reported phase transitions (from liquid to hexatic and hexatic to crystal) occurring at well-defined combinations of packing fraction and external-field magnitude. The detailed characterization of the colloidal phases, which was also performed in quasi-two-dimensional experiments of colloidal ‘hard spheres’ [64], calls for the determination of the exact functional decay of the dynamical orientational correlations shown in Fig. 3(b). These measurements would required time- or length scales that are much larger than what we can currently achieve with our experimental setup. Here, our focus is on the influence of colloidal background, as it becomes more solid-like, on the active motion of self-propelling particles. Nonetheless, the curves in Fig. 3(b) are in qualitative agreement with the existing literature and highlight the possibility of freezing or melting colloidal monolayers at fixed $\phi$ using external $\rm kHz$ AC electric fields.
@@
We then demonstrated that the tunable colloidal environments dramatically alter the trajectories of active colloids actuated by the same electric fields. In particular, during the quenching of the surrounding monolayers, the persistence time $\tau$ of the active Janus particles becomes dependent on their swimming velocity $v$. It rapidly drops by more than one order of magnitude as $v$ increases. Such decrease of persistence time (often denoted as ‘rotational diffusion enhancement’) has been observed in experiments of active colloids in glassy environments [39] and viscoelastic polymer solutions  [22, 65]. In such cases, the effect has been attributed to frictional contacts or the memory of the viscoelastic bath. As compared to these works, the combination of numerical simulations, theory and particle-resolved experiments allows to single out possible microscopic reasons behind the strong coupling between active motion and mechanical characteristics of the environments. Our results suggest that local compressions caused by the swimming force and interaction asymmetries play a major role in the anomalous reorientation. Friction and memory terms are not needed to recover the helical motion and/or rotational diffusion enhancement. As a matter of fact, our “dog chasing its tail” mechanism occurs regardless of the phase, ordering, or packing fraction of the colloidal medium, provided that the monolayer is sufficiently dense and caging effects are minimal. Under these conditions, a simple relationship between swimming velocity $v$ and persistence time $\tau$ is reported: $\tau\propto v^{-1}$. Above a critical velocity, the polar symmetry of the particles is broken and the active particle undergoes helical motion, i.e. it behaves as a chiral swimmer. This chiralization is very different from the circular motion reported for L-shaped particles [66] and modular swimmers [67, 68, 69, 70] because neither the active colloids nor the bath particles are asymmetric or bind to each other. Orbiting is here sustained by local compressed region created by the active particle itself.
@@
In summary, our findings demonstrate the ability to switch the reorientation dynamics of active particles from viscous-like to viscoelastic-like, which may have significant implications for medical applications ranging from micro-surgery to targeted drug delivery. Regardless the details of our experimental system, the microscopic mechanism of reorientation is general as it only requires a compressible medium and anisotropic interactions. We therefore envisage that, beyond regulating active-particle motion, our results will shed light on how complex environments affect the swimming behavior of microorganisms typically comprising a ‘head’ and ‘tail’.
4 Materials and Methods
@@ 4.2 Sample preparation
The two-dimensional mixtures are subjected to AC electric fields of frequency $\rm f=20$ $\rm kHz$ and magnitude up to $\rm E=108$ $\rm V/mm$ in the direction normal to the conductive substrate (see sketch in Fig. 8). The applied electric field $\rm E$ polarizes the $\rm SiO_{2}$ particles and induces long-range repulsive pair interactions whose strength increases with $\rm E$. In the case of $\rm SiO_{2}$-$\rm Au$ colloids, the metallic and dielectric hemispheres are also polarized differently, which leads to local slip flows and two-dimensional active motion of the particles with the gold cap facing forward. This mechanism is known as induced charge electrophoresis (ICEP) and the swimming velocity is controlled by electric-field magnitude $\rm E$. The electric field also orients the $\rm SiO_{2}$-$\rm Au$ particles such that the axis linking the poles of the two hemispheres is parallel to the underlying substrate, which ensures that the active motion is two dimensional at all times. Importantly, the operating frequency $\rm f=20$ $\rm kHz$ is a key parameter to observe both long-range repulsion between uncoated $\rm SiO_{2}$ particles and ICEP self-propulsion of $\rm SiO_{2}$-$\rm Au$ colloids. For example, at $\rm f\sim 1$ $\rm kHz$, the electric field drives strong solvent electrodynamic flows (EHD) along the conductive substrates in the direction of the silica microspheres [54]. This effect changes the interactions between uncoated $\rm SiO_{2}$ spheres from repulsive to attractive and swaps the swimming direction of $\rm SiO_{2}$-$\rm Au$ colloids [58].
@@ 4.3 Imaging and data analysis
The sample cell is placed onto the stage of an inverted microscope (Leica DMI3000 B, Leica Microsystems GmbH, Wetzlar, Germany), where imaging is performed with $10\times$ and $20\times$ magnification in bright-field mode. The microscope is coupled to a CMOS camera (Basler ACE) of resolution 1280 $\times$ 1080 pixels and videos are recorded at 2 frames per second. The centres of the particles are detected over time using python tracking codes (http://soft-matter.github.io/trackpy/v0.5.0/). We record approximately $15$ active trajectories for any applied electric field and packing fraction, and calculate the instantaneous velocity $\dot{\textbf{x}}$ of each particle as the distance travelled in the time interval $\rm\Delta t=2$ $\rm s$. From the magnitude and direction of $\dot{\textbf{x}}$ we compute the mean swimming velocity $v$ (or $v_{0}$ at $\phi=0$) and persistence time $\tau$ (or $\tau_{0}$ at $\phi=0$), as presented in Section 2.2. Specifically, to calculate $v$ and $v_{0}$, we first average the instantaneous velocities of each active particle $i$ to obtain its mean swimming velocity $v_{i}$ and then take the mean among all the $\sim 15$ particles. To calculate $\tau$ and $\tau_{0}$, we consider the time autocorrelation function of the instantaneous velocity vector $\rm\left\langle\dot{\mathbf{x}}_{i}(\Delta t)\cdot\dot{\mathbf{x}}_{i}(0)\right\rangle$ of the active particle $i$ [17], which decays as
@@ 4.4 Numerical simulations
4.4.1 Passive particles dynamics

### v2

@@ 2.2 Active particles in tunable colloidal environments
According to our description, $\tau\propto v^{-1}$ since more distance travelled means more reorientation events. We verified this dependence by performing numerical simulations where $v$ (swimming velocity of the individual active particles) and $\rm K$ (strength of the dipolar interactions, see Eq. (9) and Table S1) are decoupled. The results are shown in Fig. S6; in most instances, $\tau$ and $v$ are inversely proportional suggesting that each reorientation event depends weakly on the local interaction strength $\rm K$. However, since a given reorientation is amplified, the persistence time $\tau$ strongly depends on the global stiffness of the medium. The model also predicts that the magnitude of the torque acting on the Janus particle increases with $\rm E^{2}$ since this is the scaling of the dipolar repulsive interactions. Extracting the average torque as $\rm\langle M\rangle=\xi_{\rm R}\langle\dot{\theta}\rangle$ ($\langle\dot{\theta}\rangle$ is the mean angular speed, see also Eq. (10)b) reveals good agreement with $\rm\langle M\rangle\propto E^{2}$ (see black line in Fig. 4(e)) in both the numerical (Fig. 4(e), empty circles) and experimental data (Fig. 4(e), solid squares), proving that that interaction asymmetry and long-range repulsions are key ingredients to achieve self-sustaining reorientations. In a matrix of passive particles interacting only via WCA forces and in the absence of caging effects $\tau$ remains approximately constant (see Fig. S7). Likewise, no self-sustaining reorientations occur if we preserve the long-ranged dipolar interactions, but set $\boldsymbol{\ell}=0$, since the orientational and translation dynamics are decoupled.
@@ 2.3 Spontaneous chiralization
Figure 4: Self-sustaining reorientation. (a–b) Sketch of the reorientation mechanism, as modeled in numerical simulations, causing the drop of the persistence time of the active particle in colloidal environments. (a) An active particle moving with the Au cap heading produces a swimming force $f_{\rm s}\hat{\mathbf{n}}$ in the same direction. (b) Upon rotation by an angle $\Delta\theta$, the colloidal environment is compressed, generating a torque due to dipolar repulsion between passive particles and the silica back hemisphere of the active colloid. (c–d) Voronoi tessellation of the environment (c) before and (d) after a rotation. The color code indicates the local packing, i.e., the inverse area of the Voronoi cell. (e) Mean torque experienced by an active Janus particle swimming throughout a monolayer at area fraction $\phi=0.48$ for different applied electric fields. Solid and empty symbols correspond to experiments and simulations, respectively. The red dashed line connects the numerical data, and the black solid line marks the linear relationship between $\rm\langle M\rangle$ and $\rm E^{2}$. (f) Persistence time $\tau$ as a function of the average velocity $v$ in colloidal environments at different $\phi$, as indicated in the legend. The filled symbols correspond to the experiments, while the empty symbols are from numerical simulations matching the experimental values of $v$. (g) Helical active trajectory of 44-minutes duration observed in experiments when the area fraction of the surrounding colloidal bath is $\phi=0.64$. The scale bar corresponds to $20\,\mu\mathrm{m}$. (h) Numerical trajectory of an active particle swimming at velocity $v=2.4\,\mu\mathrm{m/s}$ in colloidal environments at $\phi=0.35$. The color indicates the time, from $t=0$ (light blue) to $t=40$ s (blue). The lever arm $\ell=0.4\,\mu\mathrm{m}$ determining the torque is shown as a blue arrow pointing from the geometric particle center to the interaction center. The scale bar corresponds to $1\,\mu\mathrm{m}$.
@@
The scenario depicted so far also occurs over a broad range of packing fractions $\phi$. Figure 4(f) summarizes the dependence of $\tau$ on $v$ in colloidal environments at different packing fraction: $\rm\phi=0.15$ (green data), $\rm\phi=0.35$ (blue data) and $\rm\phi=0.48$ (red data, redrawn from Fig. 3(i)). In all instances, we report a drop of persistence time as the swimming velocity increases. Only for small $v$ and $\phi$ (first three green data points) or large $v$ and $\phi$ (last two red data points), $\tau$ remains roughly constant. The observed behavior is reproduced by numerical simulations (empty symbols and connecting dotted lines) remarkably well. Note, however, that the simple model described in Fig. 4(a-d) breaks down if the environment is too sparse (at $\phi=0.15$ for small $\rm K$ values) and for strong velocity fluctuations, i.e. upon caging (at $\phi=0.48$ and $\phi=0.64$, for large $\rm K$ values). Finally, in experiments at $\phi=0.64$ a significant number of active particles undergo helical motion (Fig. 4(g) and Supplementary Video S8). They behave as if they were chiral, with orbit’s radii as large as $\sim 5\rm R_{p}$, but no preferential kick-off direction. Because both the translational and angular velocity grow linearly with $\rm E^{2}$ (Fig. 3(a) and Fig. 4(e)), the radius of the orbit, $v/\langle\dot{\theta}\rangle$, does not depend on $\rm E$. The same break of polar symmetry is observed in numerical simulations but, at $\phi=0.64$, the spontaneous chiralization is hindered by strong caging effects which, as discussed above, are more pronounced in simulations than experiments. To recover the same phenomenology reported in experiments, we keep the packing fraction relatively low ($\phi=0.35$) and virtually shift the point on which dipolar interactions act from the center of the $\rm SiO_{2}$ hemisphere towards the edge (see blue arrows in the sketch of Fig. 4(h)). Practically, this operation increases the lever arm $\ell$ and amplifies the torque exerted by the passive particles due to dipolar interactions, while keeping the swimming velocity and stiffness of the medium fixed. As the torque increases, the active colloid is bound to make a helical motion – a signature of chirality. As in experiments, there is no preferential direction of rotation.
@@
In agreement with the experimental and numerical results, $\langle\dot{\theta}\rangle$ increases linearly with the self-propulsion speed $v_{0}$, and thus vanishes in equilibrium or for a force that is applied to the geometric center of the particle ($\ell=0$). In addition, this term disappears when the local density change of the environment around the active particle is negligible, as in experiments with a low-density background, or when the background is incompressible, e.g. in a colloidal glass made of hard spheres. The full coarse-grained theory is derived and reported in Sec. 6.
@@ 3 Discussion and Conclusions
After presenting an experimental strategy to tune the structure of colloidal monolayers at fixed packing fraction, we investigated the dynamical properties of active colloids cruising throughout them. The active motion is strongly affected by the environment, which alters the particle speed and reorientation, and even gives rise to unexpected helical motion (a qualitative summary of the observed types of motion is in Fig. S8). The mechanical properties of the environment and microswimmers activity are adjusted reversibly and in situ using the same experimental control knob: the magnitude of an applied AC electric field.
@@
The role of orthogonal AC electric fields in determining the pair interaction between charged microparticles sitting onto planar electrodes has been extensively investigated in the past decades [56, 57, 58, 59]. In particular, the AC frequency is paramount since it regulates whether the interactions between particle pairs are attractive [60, 61, 62, 63] or repulsive [59, 46, 64]. At our operating frequency $\rm f=20$ $\rm kHz$, long-range dipole-dipole repulsive interactions between $\rm SiO_{2}$ particles are dominant; their strength ($\rm K$, Eq. (9)) is always much larger than the energy scale of the WCA potential ($\epsilon$, Eq. (8)) at the typical experimental inter-particle distances identified by the first peak of the pair correlation function (Fig. 2(a), Fig. S1 and Table S1). As such, the experimental system presented in Fig. 2 is similar to the colloidal monolayers described in Refs. [59, 65, 66], where quenching (or melting) took place under external electric ($\rm f>1$ $\rm MHz$) and magnetic fields. These works reported phase transitions (from liquid to hexatic and hexatic to crystal) occurring at well-defined combinations of packing fraction and external-field magnitude. The detailed characterization of the colloidal phases, which was also performed in quasi-two-dimensional experiments of colloidal ‘hard spheres’ [67], calls for the determination of the exact functional decay of the dynamical orientational correlations shown in Fig. 2(g). These measurements would required times or length scales that are much larger than what we can currently achieve with our experimental setup. Here, our focus is on the influence of colloidal background, as it becomes more quenched, on the active motion of self-propelling particles. Nonetheless, the curves in Fig. 2(g) are in qualitative agreement with the existing literature and highlight the possibility of freezing (or melting) colloidal monolayers at fixed $\phi$ using external $\rm kHz$ AC electric fields.
@@
We then demonstrated that the tunable colloidal environments dramatically alter the trajectories of active colloids actuated by the same electric fields. In particular, during the quenching of the surrounding monolayers, the persistence time $\tau$ of the active Janus particles becomes dependent on their swimming velocity $v$. It rapidly drops by more than one order of magnitude as $v$ increases. A similar decrease of persistence time (often denoted as ‘rotational diffusion enhancement’) has been reported for self-propelling Janus particles in polymer solutions  [25, 24, 22] and initially ascribed to viscoelasticity and memory effects. The same experimental results were recently revised by De Corato et al. using a memory-free model based on local density variations, anisotropic interactions and advection terms [27, 26]. The first two ingredients are also present in our “dog chasing its tail”, but the advection of background particles is absent since hydrodynamic interactions are neglected. Local density fluctuations are here rather due to local compressions caused by the swimming force of the Janus particle. They occur regardless of the phase, ordering, or packing fraction of the colloidal medium, provided that the monolayer is sufficiently dense and caging effects are minimal. Under these conditions, a simple relationship between swimming velocity $v$ and persistence time $\tau$ is reported: $\tau\propto v^{-1}$. Above a critical velocity, the polar symmetry of the particles is broken and the active particle undergoes helical motion, i.e. it behaves as a chiral swimmer. This chiralization is very different from the circular motion reported for L-shaped particles [68] and modular swimmers [69, 70, 71, 72] because neither the active colloids nor the bath particles are asymmetric or bind to each other. Instead, orbiting is sustained by local compressed regions created by the active particle itself.
@@
In summary, our findings demonstrate the ability to tune in situ the reorientation dynamics of active particles, which may have significant implications for medical applications ranging from micro-surgery to targeted drug delivery. Regardless the details of our experimental system, the microscopic mechanism of reorientation is general as it only requires a compressible medium and anisotropic interactions. We therefore envisage that, beyond regulating active-particle motion, our results will shed light on how compressible micro- or nanostructured fluids affect the swimming behavior of microorganisms typically comprising a ‘head’ and ‘tail’. For example, catalytic micromotors, which are Janus just like our active particles, may exhibit a similar motion when added to muddy waters for wastewater treatments or environmental monitoring. Similarly, the body and flagella of bacteria may absorb/repel macrosolutes (e.g. polymers) in a different manner leading to local compressed/depleted regions in the environmental fluid.
4 Experimental Section
@@ 4.2 Sample preparation
The two-dimensional mixtures are subjected to AC electric fields of frequency $\rm f=20$ $\rm kHz$ and magnitude up to $\rm E=108$ $\rm V/mm$ in the direction normal to the conductive substrate (see sketch in Fig. 5). The applied electric field $\rm E$ polarizes the $\rm SiO_{2}$ particles and induces long-range repulsive pair interactions whose strength increases with $\rm E$. In the case of $\rm SiO_{2}$-$\rm Au$ colloids, the metallic and dielectric hemispheres are also polarized differently, which leads to local slip flows and two-dimensional active motion of the particles with the gold cap facing forward. This mechanism is known as induced charge electrophoresis (ICEP) and the swimming velocity is controlled by electric-field magnitude $\rm E$. The electric field also orients the $\rm SiO_{2}$-$\rm Au$ particles such that the axis linking the poles of the two hemispheres is parallel to the underlying substrate, which ensures that the active motion is two dimensional at all times. Importantly, the operating frequency $\rm f=20$ $\rm kHz$ is a key parameter to observe both long-range repulsion between uncoated $\rm SiO_{2}$ particles and ICEP self-propulsion of $\rm SiO_{2}$-$\rm Au$ colloids. For example, at $\rm f\sim 1$ $\rm kHz$, the electric field drives strong solvent electrodynamic flows (EHD) along the conductive substrates in the direction of the silica microspheres [57]. This effect changes the interactions between uncoated $\rm SiO_{2}$ spheres from repulsive to attractive and swaps the swimming direction of the $\rm SiO_{2}$-$\rm Au$ colloids [61].
@@ 4.3 Imaging and data analysis
The sample cell is placed onto the stage of an inverted microscope (Leica DMI3000 B, Leica Microsystems GmbH, Wetzlar, Germany), where imaging is performed with $10\times$ and $20\times$ magnification in bright-field mode. The microscope is coupled to a CMOS camera (Basler ACE) of resolution 1280 $\times$ 1080 pixels and videos are recorded at 2 frames per second. The centres of the particles are detected over time using python tracking codes (http://soft-matter.github.io/trackpy/v0.5.0/). We record approximately $15$ active trajectories for any applied electric field and packing fraction, and calculate the instantaneous velocity $\dot{\textbf{x}}$ of each particle as the distance travelled in the time interval $\rm\Delta t=2$ $\rm s$. From the magnitude and direction of $\dot{\textbf{x}}$ we compute the mean swimming velocity $v$ (or $v_{0}$ at $\phi=0$) and persistence time $\tau$ (or $\tau_{0}$ at $\phi=0$). Specifically, to calculate $v$ and $v_{0}$, we first average the instantaneous velocities of each active particle $i$ to obtain its mean swimming velocity $v_{i}$ and then take the mean among all the $\sim 15$ particles. To calculate $\tau$ and $\tau_{0}$, we consider the time autocorrelation function of the instantaneous velocity vector $\rm\left\langle\dot{\mathbf{x}}_{i}(\Delta t)\cdot\dot{\mathbf{x}}_{i}(0)\right\rangle$ of the active particle $i$ [17], which decays as
@@ 5 Numerical Section
5.1 Passive particles dynamics